Roscoe and Bruxie can both produce Chicken and Waffles. Roscoe’s is in LA county and Bruxie’s is in Orange county. It takes Roscoe 5 minutes to make a waffle and 20 minutes to prepare chicken. It takes Bruxie 10 minutes to make a waffle and 20 minutes to prepare chicken. What is Roscoe's opportunity cost for making a waffle?

a.

1/4 Chicken

b.

4 Chickens

c.

1/3 Chicken

d.

1/4 Waffle

e.

1/3 Waffle

f.

1/8 Chicken
Answer #1

Ans) the correct option is a) 1/4 chicken

Roscoe's opportunity cost for making a waffle = 5/20 chicken = 1/4 chicken